how do you find the vertex of the parabola for the quadratic equation
y=x^2-2x-35

Answer by scott8148(6628)

completing the square ___ y = x^2 - 2x + 2 - 35 - 2

y = (x - 1)^2 - 37 ___ this is the vertex form , y = a(x - h)^2 + k ; the vertex is (h,k)